# Normal Quaternionic Matrices and Finitely Generated Witt Rings

### v1

@@ 6. The Computation
    Among the parameters for $n=7,s=1$, there was one parameter which was handled in about 7 minutes, one in about 3 minutes and all the others in less than a minute. For most of the parameters, the computations were done in only a few seconds.
  5. (e)
@@
    • From a purely computational point of view, it does not seem to be obvious a priori which parameters (or subcases thereof) need the most runtime. It appears that the most time consuming parameters usually have large $d$ and small values for $v-d$ but some parameters that satisfy these conditions could be handled with surprising ease. Additionally, while the ’critical’ parameters usually satisfy $r=0$, there was one exception for order $7$ and level $4$, namely $(4,2,1,4,1,1)$, which proved to be extremely difficult to handle. Criteria to identify these critical parameters and their most time consuming subcases would enable us to apply automated strategies for parallelization of calculations, instead of relying on manual inspections for each case.
@@
As mentioned before, Carson and Marshall first computationally verified the ETC for order $5$ in their 1982 article [CarsonMarshall_Decomposition]. To compute all abstract Witt rings for a given group $G$ of exponent $2$, they consider the value groups $V(a)$ for $a\in G$. There are some similarities to the approach presented here, as these value groups can be recovered by considering the zeroes in the rows of the completion $\overline{M}$ of a given quaternionic matrix in the following way: if $a\in G$ is associated with row and column $k_{a}$ in the completion, the set $V(a)$ consists of all $b\in G$ such that $\overline{M}_{k_{a},k_{b}}=0$ for $b$ associated with row and column $k_{b}$.
@@ Appendix A Lists of Normal Quaternionic Matrices
We include the list of all normal quaternionic matrices of all quaternionic structures of order $4$ and all those of orders $5$ and $6$ that are nondegenerate, nonpythagorean and nonrigid. Instead of displaying the entire matrices, we only include the upper right triangle (row by row, including the first diagonal entry) as this uniquely describes the quaternionic matrix (along with the level). For the same level, the matrices are sorted lexicographically.
normal quaternionic matrices of order 4
level 1:
$\mathbb{L}_{1,0}^{4}$ (0,0,0,0,0,0,0)
$\mathbb{L}_{1,0}^{2}\times\mathbb{L}_{1,0}\Delta$ (0,0,0,0,0,0,1)
$\mathbb{L}_{1,0}\times\mathbb{L}_{1,0}^{2}\Delta$ (0,0,0,0,0,1,2)
$\mathbb{L}_{1,0}\times\mathbb{L}_{1,0}\Delta_{2}$ (0,0,0,0,1,2,4)
$\mathbb{L}_{1,0}^{3}\Delta$ (0,0,0,1,0,2,4)
$\mathbb{L}_{4,0}$ (0,0,0,1,1,0,0)
$\mathbb{L}_{1,0}\Delta^{2}$ (0,0,0,1,2,0,0)
$(\mathbb{L}_{1,0}\times\mathbb{L}_{1,0}\Delta)\Delta$ (0,0,0,1,2,4,8)
$\mathbb{L}_{1,0}^{2}\Delta_{2}$ (0,0,1,2,4,8,16)
$\mathbb{L}_{1,0}\Delta_{3}$ (0,1,2,4,8,16,32)
level 2:
$\mathbb{L}_{1,1}^{4}$ (0,0,0,0,0,0,0)
$\mathbb{L}_{1,1}^{2}\times\mathbb{L}_{1,0}\Delta$ (0,0,0,0,0,0,1)
$\mathbb{L}_{1,1}\times\mathbb{L}_{1,0}^{2}\Delta$ (0,0,0,0,0,1,2)
$\mathbb{L}_{1,1}\times\mathbb{L}_{1,0}\Delta_{2}$ (0,0,0,0,1,2,4)
$\mathbb{L}_{1,1}^{2}\times\mathbb{L}_{1,1}\Delta$ (0,0,0,1,0,0,0)
$\mathbb{L}_{1,1}\times\mathbb{L}_{1,1}^{2}\Delta$ (0,0,0,1,0,0,2)
$\mathbb{L}_{1,1}^{3}\Delta$ (0,0,0,1,0,2,4)
$\mathbb{L}_{4,1}$ (0,0,0,1,1,0,0)
$\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}\Delta$ (0,0,0,1,2,0,0)
$(\mathbb{L}_{1,1}\times\mathbb{L}_{1,0}\Delta)\Delta$ (0,0,0,1,2,4,8)
$\mathbb{L}_{1,1}\times\mathbb{L}_{1,1}\Delta_{2}$ (0,0,1,2,0,0,4)
$\mathbb{L}_{1,1}\Delta^{2}$ (0,0,1,2,0,2,0)
$(\mathbb{L}_{1,1}\!\!\times\!\!\mathbb{L}_{1,1}\Delta)\Delta$ (0,0,1,2,0,4,8)
$\mathbb{L}_{1,1}^{2}\Delta_{2}$ (0,0,1,2,4,8,16)
$\mathbb{L}_{1,1}\Delta_{3}$ (0,1,2,4,8,16,32)
level $>2$:
$\mathbb{L}_{1,1}^{3}\times\mathbb{L}_{1}$ (1,0,0,0,0,0,0)
$\mathbb{L}_{1,1}^{2}\times\mathbb{L}_{1}^{2}$ (1,0,0,2,0,0,0)
$\mathbb{L}_{1,1}\times\mathbb{L}_{1}^{3}$ (1,0,2,4,0,0,0)
$\mathbb{L}_{1,1}\times\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta$ (1,0,0,0,0,0,2)
$\mathbb{L}_{1,1}\times\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta$ (1,0,0,2,0,0,2)
$\mathbb{L}_{1,1}\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta$ (1,0,0,2,0,0,4)
$\mathbb{L}_{1,1}\times\mathbb{L}_{1}^{2}\Delta$ (1,0,2,4,0,0,8)
$\mathbb{L}_{1,1}\times\mathbb{L}_{3}$ (1,0,0,0,0,0,1)
$\mathbb{L}_{1}^{4}$ (1,2,4,8,0,0,0)
$\mathbb{L}_{1}^{2}\times\mathbb{L}_{1,0}\Delta$ (1,0,0,2,4,0,0)
$\mathbb{L}_{1}^{2}\times\mathbb{L}_{1,1}\Delta$ (1,0,2,4,0,4,0)
$\mathbb{L}_{1}\times\mathbb{L}_{1,0}^{2}\Delta$ (1,0,0,0,0,2,4)
$\mathbb{L}_{1}\times\mathbb{L}_{1,1}^{2}\Delta$ (1,0,0,2,0,2,4)
$\mathbb{L}_{1}\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta$ (1,0,2,4,0,4,8)
$\mathbb{L}_{1}\times\mathbb{L}_{1}^{2}\Delta$ (1,2,4,8,0,0,16)
$\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta_{2}$ (1,0,0,0,2,4,8)
$\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta_{2}$ (1,0,2,4,2,4,8)
$\mathbb{L}_{1}\times\mathbb{L}_{3}$ (1,0,0,2,2,0,0)
$(\mathbb{L}_{1,1}^{2}\times\mathbb{L}_{1})\Delta$ (1,0,0,2,0,4,8)
$(\mathbb{L}_{1,1}\times\mathbb{L}_{1}^{2}$ (1,0,2,4,0,8,16)
$\mathbb{L}_{1}^{3}\Delta$ (1,2,4,8,0,16,32)
$(\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta)\Delta$ (1,0,0,2,4,8,16)
$(\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta)\Delta$ (1,0,2,4,2,8,16)
$(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta_{2}$ (1,0,2,4,8,16,32)
$\mathbb{L}_{1}^{2}\Delta_{2}$ (1,2,4,8,16,32,64)
$\mathbb{L}_{3}\Delta$ (1,0,0,2,1,4,8)
normal quaternionic matrices of order 5 (nondegenerate, nonpythagorean, nonrigid) level 1: $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,0}^{2}\Delta$ (0,0,0,0,1,0,0,2,4,0,0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,0}\Delta_{2}$ (0,0,0,0,1,2,4,0,8,0,0) level 2: $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}^{2}\Delta$ (0,0,0,0,1,0,0,2,4,0,0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,0}^{2}\Delta$ (0,0,0,0,1,0,2,0,4,0,0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,0}\Delta_{2}$ (0,0,0,0,1,2,4,0,8,0,0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,1}^{2}\Delta$ (0,0,0,1,2,0,0,2,0,4,0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}\Delta_{2}$ (0,0,0,1,2,4,0,0,0,0,8) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,1}\Delta_{2}$ (0,0,1,2,4,0,0,4,8,0,0) level $>2$: $\mathbb{L}_{5}$ (1,0,0,0,0,0,0,1,1,0,0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{3}$ (1,0,0,0,0,0,0,1,2,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{4,0}$ (1,0,0,0,0,0,0,2,2,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta^{2}$ (1,0,0,0,0,0,0,2,4,0,0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{3}$ (1,0,0,0,2,0,0,2,1,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{4,1}$ (1,0,0,0,2,0,0,2,2,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}\Delta$ (1,0,0,0,2,0,0,2,4,0,0) $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta$ (1,0,0,0,2,0,0,4,8,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta^{2}$ (1,0,0,2,4,0,0,4,2,0,0) $\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta$ (1,0,0,2,4,0,0,4,8,0,0) $\mathbb{L}_{1}^{2}\times\mathbb{L}_{3}$ (1,0,0,2,4,2,0,0,0,0,0)
@@
(nondegenerate, nonpythagorean, nonrigid) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,0}^{3}\Delta$ (0,0,0,0,0,1,0,0,0,2,0,0,4,8,0,0) $\mathbb{L}_{1,0}^{2}\Delta^{2}$ (0,0,0,0,0,1,0,0,0,2,0,4,0,8,0,0) $\mathbb{L}_{1,0}^{2}\Delta\times\mathbb{L}_{1,0}\Delta_{2}$ (0,0,0,0,0,1,0,0,0,2,4,8,0,16,0,0) $\mathbb{L}_{6,0}$ (0,0,0,0,0,1,0,0,1,0,1,0,0,0,0,0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{4,0}$ (0,0,0,0,0,1,0,0,1,0,2,0,0,0,0,0) $\mathbb{L}_{1,0}\Delta^{3}$ (0,0,0,0,0,1,0,0,2,0,4,0,0,0,0,0) $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1,0}\times\mathbb{L}_{1,0}\Delta)\Delta$ (0,0,0,0,0,1,0,0,2,0,4,0,8,0,16,0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,0}^{2}\Delta_{2}$ (0,0,0,0,0,1,0,2,4,0,8,16,0,32,0,0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,0}\Delta_{3}$ (0,0,0,0,0,1,2,4,8,0,16,32,0,64,0,0) $\mathbb{L}_{1,0}\Delta_{2}^{2}$ (0,0,0,0,1,2,4,8,0,0,16,0,0,0,0,32)

### v2

@@ 4 $\units0\,\unit${s} $\units0\,\unit${s} $\units12\,\unit${s} $204\,\text{days}$
We now point out the parallelization strategies used to deal with the case $n=7$. In a first step, for each choice of $s\in\{1,2,4\}$, we applied the algorithm parallelized for the different choices of parameters $s,d,f,r,t,u,v$, i.e. the different choices of the first two rows of the normal quaternionic matrices, as presented in Proposition 5.5. For level $s=1$, there were $41$ parameters to consider, the most expensive one of them having a computation time of roughly $7$ minutes.
For level $s=2$ resp. $s\geq 4$, the number of parameters to consider is $72$ resp. $196$. Among these, there were $3$ resp. $12$ with a computation time of at least $\unith5\,\unit${h} whose isolated runtimes (again assumed to have run on a single core) can be found in Table 4. This illustrates how most of the computation time focuses on a very small number of parameters.
parameter runtime
$s=2$:
$(4,1,1,4,1,1)$ 22 days, $\unith11\unit${h}
$(4,1,0,4,1,1)$ $\unith15\unit${h}
$(4,1,0,4,2,2)$ $\unith11\unit${h}
remaining $69$ parameters $\unith2\unit${h}
$s\geq 4$:
$(4,1,0,4,1,1)$ $85$ days, $\unith5\unit${h}
$(4,2,1,4,1,1)$ $71$ days, $\unith16\unit${h}
$(4,1,0,4,2,2)$ $40$ days, $\unith7\unit${h}
$(4,1,0,4,2,3)$ 11 day, $\unith15\unit${h}
$(3,1,0,3,0,0)$ 11 day, $\unith9\unit${h}
$(3,1,0,3,1,1)$ $\unith21\unit${h}
$(4,1,0,4,3,3)$ $\unith10\unit${h}
$(3,1,0,3,0,1)$ $\unith9\unit${h}
$(4,1,0,4,0,1)$ $\unith9\unit${h}
$(4,1,0,4,0,0)$ $\unith8\unit${h}
$(2,1,0,3,0,0)$ $\unith7\unit${h}
$(4,1,0,4,1,2)$ $\unith5\unit${h}
remaining $184$ parameters $\unith19\unit${h}
Table 4. parameters with runtimes of at least $\unith5\unit${h} for $n=7$ and $s=2$, $s\geq 4$
For these parameters, we applied further parallelization techniques which will be explained in detail in Example 6.7.
Example 6.7.
We use the case $n=7$, $s\geq 4$ with parameters $(4,1,0,4,1,1)$ to illustrate our parallelization strategies for the most time consuming parameters. This strategy was applied for all parameters with runtimes of more than $24$ hours and - during the process of adapting the algorithm - also for some of the other parameters and for which we also established unparallelized results.
In the implementation of Algorithm 1, we added the option to have a list of partial quaternionic matrices (with all their class attributes) as an input from which we can resume the computation to complete the given partial matrices. Further, built up from auxiliary functions used for debugging, we implemented the option to only compute all quaternionic matrices for a given input (either a set of parameters of a list of incomplete quaternionic matrices as just explained) just up to a certain entry $(i,j)$ specified in the code. The advantage of this is that it is possible to partition the case into many sub cases which can be computed on different kernels as they have no interdependence. In practice, it was possible for us to compute up to approx. 60000 cases in one run, resulting in .obj-files of about 3 GB, serving as the input for a future run.
For the example under consideration, we get the first two rows of the normal quaternionic matrix as follows:
$\begin{pmatrix}1&0&0&2&4&8&16\\ 0&0&0&0&32&64&128\end{pmatrix}$
In a first step, $16$ partial quaternionic matrices up to entry $(2,3)$ which is the first entry not determined by the symmetry and the diagonal conditions were calculated. While $11$ of those cases could be computed entirely in a few seconds, the other $5$ cases, namely
$\begin{pmatrix}1&0&0&2&4&8&16\\ 0&0&0&0&32&64&128\\ 0&0&0&x&&&\end{pmatrix}$
for $x\in\{0,2,33,40,256\}$, were partially completed up to a point to reach a number of cases that is high enough to ensure efficient parallelization but without exceeding sensible limits for storage.
Once again, we only illustrate the most complicated case which proved to be $x=256$ above; some discussions on which (sub) cases are the most complex and how it might be possible to reduce computation times with more theoretic input can be found in the following Remark 6.8.
The algorithm computed $41172$ cases up to entry $(3,4)$. All but $10$ of those were computed without any further intermediate steps, most of them with case times of under a minute, but some with case times of a few hours as well. The remaining $10$ cases were again only partially computed up to entry $(3,5)$ or $(4,4)$, each dividing into $300$ to $60000$ subcases which could be all be computed in a few hours at most.
For each case, a result file was produced, including the runtime of the case, the partial quaternionic matric used as the input, and the number of results of quaternionic matrices. In the given example, a simple routine checked all the result files to add the number of quaternionic matrices computed and also the different case times. As conjectured in the ETC, no quaternionic matrices were computed while the computation times for the parameter considered added up to $85$ days and $5$ hours, see Table 4.
Remark 6.8.
@@
  • From a purely computational point of view, it does not seem to be obvious a priori which parameters (or subcases thereof) need the most runtime. It appears that the most time consuming parameters usually have large $d$ and small values for vdv-d but some parameters that satisfy these conditions could be handled with surprising ease. Additionally, while the ’critical’ parameters usually satisfy $r=0$, there was one exception for order $7$ and level $4$, namely $(4,2,1,4,1,1)$, which proved to be the second most time consuming parameter. Criteria to identify these critical parameters and their most time consuming subcases would enable us to apply automated strategies for parallelization of calculations, instead of relying on manual inspections for each case.
  • Aside from the limitations due to the common slot property, the only condition that enables us to exclude new powers of $2$ as the next entry comes from excluding rigid and in some cases $2$-rigid elements. Some test runs indicate that excluding $2$-rigid elements entirely would massively reduce the computation time for $n=7$, $s\geq 4$ to about $1$ day. Hence, it would be of great help to generalize the results obtained by Carson and Marshall resp. Cordes on $2$-rigid elements, which are mentioned in Proposition 2.8.
@@
As mentioned before, Carson and Marshall first computationally verified the ETC for order $5$ in their 1982 article [CM82]. To compute all abstract Witt rings for a given group $G$ of exponent $2$, they consider the value groups $V(a)$ for $a\in G$. There are some similarities to the approach presented here, as these value groups can be recovered by considering the zeros in the rows of the completion $\overline{M}$ of a given quaternionic matrix in the following way: if $a\in G$ is associated with row and column $k_{a}$ in the completion, the set $V(a)$ consists of all $b\in G$ such that $\overline{M}_{k_{a},k_{b}}=0$ for $b$ associated with row and column $k_{b}$.
@@
Acknowledgments
Significant parts of the calculations for this publication were performed on the HPC cluster Elysium of the Ruhr University Bochum, subsidised by the DFG (INST 213/1055-1).
@@ Appendix A Lists of Normal Quaternionic Matrices
We include the list of all normal quaternionic matrices of all quaternionic structures of order $6$ that are nondegenerate, nonpythagorean and nonrigid. Instead of displaying the entire matrices, we only include the upper right triangle (excluding the diagonal) as this uniquely describes the quaternionic matrix (along with the level) and indicate line breaks by a semicolon. For the same level, the matrices are sorted lexicographically. Those matrices that do not fulfill the additional rigidity assumptions in Proposition 2.8 and thus not computed with our algorithm are marked with an asterisk.
@@
(nondegenerate, nonpythagorean, nonrigid) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,0}^{3}\Delta$ (0,0,0,0,1; 0,0,0,2; 0,0,4; 8,0; 0) $\mathbb{L}_{1,0}^{2}\Delta^{2}$ (0,0,0,0,1; 0,0,0,2; 0,4,0; 8,0; 0) $\mathbb{L}_{1,0}^{2}\Delta\times\mathbb{L}_{1,0}\Delta_{2}$ (0,0,0,0,1; 0,0,0,2; 4,8,0; 16,0; 0) $\mathbb{L}_{6,0}$ (0,0,0,0,1; 0,0,1,0; 1,0,0; 0,0; 0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{4,0}$ (0,0,0,0,1; 0,0,1,0; 2,0,0; 0,0; 0) $\mathbb{L}_{1,0}\Delta^{3}$ (0,0,0,0,1; 0,0,2,0; 4,0,0; 0,0; 0) $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1,0}\times\mathbb{L}_{1,0}\Delta)\Delta$ (0,0,0,0,1; 0,0,2,0; 4,0,8; 0,16; 0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,0}^{2}\Delta_{2}$ (0,0,0,0,1; 0,2,4,0; 8,16,0; 32,0; 0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,0}\Delta_{3}$ (0,0,0,0,1; 2,4,8,0; 16,32,0; 64,0; 0) $\mathbb{L}_{1,0}\Delta_{2}^{2}$ (0,0,0,1,2; 4,8,0,0; 16,0,0; 0,0; 32)